Mitchell paid $94 for ten tickets to the concert. Adult tickets cost $15 each. Students tickets cost $8 each. How many of each kind of tickets did he buy?
Answer by mananth(16949)

x Number of adult tickets
y Number of student tickets
Total attended
7 x + 10 y = 10 .............1

15 x + 8 y = 94 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 5
-26 x -40 y = -40
75 x + 40 y = 470
Add the two equations
49 x = 430
/ 49
x = 9
plug value of x in (1)
7 x + 10 y = 10
57 + 10 y = 10
10 y = 10 -57
10 y = -47
y = -5
$ 9 Number of adult tickets
$ -5 Number of student tickets
m.ananth@hotmail.ca